Two negatively charged particles of unit magnitude and charge are placed along a straight line. If charge q is placed at the midpoint between charges, then, q is in
a) stable equilibrium
b) unstable equilibrium
c) neutral equilibrium
d) none

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
see
assume distance between the particles -1c ans -1c is 2d
then q charge is at d
assume q is positve (without loss of generality)
then attr for by left -1c =kq/d^2
attr. right -1c =kq=/d^2
lets assume particle suffers a small diplacement of x x<<d towards right particle

now force left= kq/(d+x)^2
right =kq(/d-x)^2
since x<<d

we approximate the force and find difference right -left =4kqx/d^2 (1+x)^n=1+nx when x<<1
which is not proportional to -x
so unstable equilibirum